OpenStudy (anonymous):

The material used in nuclear bombs is Pu 239 with a half-life of about 20,000 years. How long must we wait for a buried stockpile of this substance to decay to 4% of its original Pu 239 mass? I got t = 92896.85 years approx= 93,000 years

OpenStudy (anonymous):

The half-life equation: \[y=Ce^{rt}\] You're given that the half-life is 20,000, which means that at \(t=20,000\), \(C\) will have decayed to half its original value. \[\frac{1}{2}C=Ce^{20,000r}\] or \[\frac{1}{2}=e^{20,000r}\] Solve for \(r\): \[\frac{\ln(1/2)}{20,000}=r\] To find the time it takes for 4% to remain, we solve for \(t\): \[0.04=e^{rt}\\ \frac{\ln0.04}{r}=t\approx 92,877\] The discrepancy between our results is probably a consequence of rounding.

OpenStudy (anonymous):

thank you.